### v1

@@ Abstract
keywords:
classical Schwarz methods, multiprecision computation, M-matrix, convergence analysis
{AMS}
65F10, 65F08, 65Y99, 65Y99
@@ 1 Introduction
(1) $A\mathbf{u}=\mathbf{f},\quad A\in\mathbb{R}^{N\times N},\mathbf{f}\in\mathbb{R}^{N}.$
@@ 2 Algebraic Schwarz methods
(2) $\bar{R}_{i}=\begin{bmatrix}I_{\bar{N}_{i}}0\end{bmatrix}\overline{\Pi}_{i}\in\mathbb{R}^{\bar{N}_{i}\times N},$
@@
(3) $\mathbf{u}^{(n+1)}=T_{\star}\mathbf{u}^{(n)}+\mathbf{c}_{\star},\quad\mathrm{or,\;equivalently,\;for\;the\;errors\;}\quad\mathbf{e}^{(n+1)}=T_{\star}\mathbf{e}^{(n)},$
where $\star\in\{\mathrm{AS},\mathrm{RAS},\mathrm{MS}\}$, $\mathbf{c}_{\star}\in\mathbb{R}^{N}$ are some constant vectors, $\mathbf{e}^{(n)}:=\mathbf{u}-\mathbf{u}^{(n)}\in\mathbb{R}^{N}$ is the error vector after $n$ iterations and the matrices $T_{\star}$ are the iteration matrices of the respective methods, given by***An equivalent multi-splitting formulation of these can be found in [18, Section 2] and [3, Sections 2 and 3].
@@
(5) $M_{\mathrm{MS}}^{-1}=\left(I_{N}-T_{\mathrm{MS}}\right)A^{-1},$
@@
(6) $A^{T}=A\quad\mathrm{and}\quad\mathbf{v}^{T}A\mathbf{v}>0\;\mathrm{for\;all\;}\mathbf{v}\neq 0.$
@@
Theorem 2.1 ([17, Lemma 2.8][3, Theorem 3.8]).
@@
Theorem 2.2 ([17, Theorem 3.4][18, Theorem 4.4][3, Theorem 3.5]).
@@
Remark 2.3.
We note that Theorem 2.2 seemingly guarantees RAS convergence even for “no overlap” case where $W_{i}=\overline{W}_{i}$ for some (or all) $i$. At first this might seem contradictory to the analytic results for the standard Laplace test problem, where $W_{i}$ corresponds to discretization of the problem on $\Omega_{i}$. But there is no contradiction as even the “no overlap” case in the algebraic sense corresponds to the “$h$ overlap” in the analytic sense of the subdomains $\Omega_{i}$, thanks to the Dirichlet boundary condition enforcement.
@@ 3 Multiprecision computations
The chop toolbox is an open-source MATLAB toolboxTowards the end of preparing the manuscript, the chop toolbox has been also released for python, see [9]. developed for simulating different precisions using the native double of MATLAB, essentially by removing a portion of the mantissa of the result after each operation, corresponding to “rounding” back to the simulated precision. For computations in single precision or lower, chop faithfully simulates the computation in the precision (see [31, Section 3.1]) and can also simulate the underflow/overflow during the computation. Although this toolbox outperforms many other options (see [31, Sections 5 and 6]), it makes some computations prohibitively time-consuming, even after adapting it to sparse matrices. Although it allows for arbitrary user-defined formats (defined by the number of bits allocated to the exponent and the significand), we will restrict ourselves to the currently standard ones, summarized in Table 1 below.
@@
(8) $\mathcal{A}_{i}\mathbf{v}_{i}=\mu\mathbf{b}_{i}$
@@
(9) $\mathcal{A}_{i}\hat{\mathbf{v}}_{i}=\hat{\mathbf{b}}_{i},$
@@ 4 Algebraic analysis of multiprecision Schwarz methods
In this section we give analogous results to Theorems 2.1 and 2.2 when the subdomain solves $A_{i}^{-1}$ are represented using a lower-precision in some way. The purpose of the numerical experiments here is twofold – to demonstrate the theoretical results and also to build an intuition for and understanding of the multiprecision Schwarz methods. Therefore, we will use the convergence properties (such as number of iterations or the convergence factor) to compare the results with their “full precision” counterparts, as opposed to, e.g., runtimes. All of the code used to produce these is available at https://github.com/MichalOutrata/mpSchwarz but, naturally, the code assumes that both the advanpix as well as the chop toolboxes are available.
@@
(11) $A_{i}=\tilde{A}_{i}-\left(\tilde{A}_{i}-A_{i}\right),\quad i=1,\dotsc,p\;.$
@@ 4.1 The general case
(12) $\tilde{A}_{i}^{-1}\geq 0\quad\mathrm{and}\quad\tilde{A}_{i}^{-1}\left(\tilde{A}_{i}-A_{i}\right)=\tilde{A}_{i}^{-1}\mathtt{E}_{i}\geq 0.$
@@
(13) $\tilde{A}_{i}=\mu^{-1}(D_{i}^{r})^{-1}\tilde{\mathcal{A}}_{i}(D_{i}^{r})^{-1}$
@@
Here we would like to recall a useful observation for diagonal re-scaling of general stationary iterative methodsWe came across this observation in [29, Section 17.2, below eqn. (17.3)] but this is likely not the original reference. – if the stationary iterative method is based on a splitting A=MNA=M-N such that the entries of $M$ are multiples of the corresponding entries of $A$, then the iteration matrices for $A$ and $D_{1}AD_{2}$ are similar (i.e., have the same convergence factor) for any non-singular diagonal matrices $D_{1},D_{2}$. On one hand, this shows that for $u_{\ell}=u_{w}$, the re-scaling above doesn’t affect the asymptotic convergence rate in our case. On the the other, we are clearly interested in cases where the rounding does make a difference and through this observation we see that we can expect the convergence factor to be affected by the re-scaling.
@@
(14) $\tilde{\mathcal{A}}_{i}^{-1}\geq 0\quad\mathrm{and}\quad\tilde{\mathcal{A}}_{i}^{-1}\left(\tilde{\mathcal{A}}_{i}-\mathcal{A}_{i}\right)=\tilde{\mathcal{A}}_{i}^{-1}\mathtt{F}_{i}\geq 0.$
@@
(15) $\tilde{\mathcal{A}}_{i}^{-1}=\mathcal{A}_{i}^{-1}\left(I+\mathtt{F}_{i}\mathcal{A}_{i}^{-1}\right)^{-1},$
@@
(17) $\tilde{\mathcal{A}}_{i}^{-1}=\mathcal{A}_{i}^{-1}\sum\limits_{k=0}^{+\infty}\left(-\mathtt{F}_{i}\mathcal{A}_{i}^{-1}\right)^{k},$
@@
(19) $\mathcal{A}_{i}^{-1}\geq\mathcal{A}_{i}^{-1}\mathtt{F}_{i}\mathcal{A}_{i}^{-1},$
@@
(20) $|\mathtt{F}_{i}|\leq u_{\ell}|\mathcal{A}_{i}|,$
which is generally guaranteed (the absolute value is to be understood component-wise). Also, similarly to [2, Sections 4 and 5], both (16) and (19) invite us to choose $(u_{\ell})_{i}$ for each subdomain independently, based on the relevant quantities or their estimates§§§In [2], the authors work with a similar idea but calculate explicitly the analogue of the inverses $\tilde{\mathcal{A}}_{i}^{-1}$ in different precisions based on their conditioning. This is somewhat complementary to our approach as our interest lies in the analysis of the resulting method rather than in the practical aspect, which has been covered in [2] and we do not comment further on how to choose $u_{\ell}$ (or $(u_{\ell})_{i}$) for the subdomain problems.. We keep (16) and (19) as assumptions, coupling the subdomain problems and the choice of the lower precision $u_{\ell}$ and move our attention to the condition $\mathtt{F}_{i}\geq 0$.
@@
To this end we assume that for any precision $u_{\ell}$ we have at our disposal the functions $\mathtt{rd}_{u_{\ell}}$ and $\mathtt{ru}_{u_{\ell}}$ that round towards zero (down) and towards plus/minus infinity (up)In the chop toolbox, these are already implemented and for the advanpix package, these are straight-forward to implement as we deal with the precision $u_{\ell}$ corresponding to $d_{\ell}$ accurate decimal digits (as opposed to dealing with bits in the case of chop).. We then introduce the rounding procedure $\mathrm{round}_{Mmtrx}()$ that for any matrix $X$ gives its low-precision approximation $\mathrm{round}_{Mmtrx}(X)$ given by
@@
(21) $\tilde{\mathcal{A}}_{i}=\mathrm{round}_{Mmtrx}(\mathcal{A}_{i}),$
we get $\mathtt{F}_{i}\geq 0$ and obtain a convergent multiprecision Schwarz methods under the assumptions (16) and  (19); we summarize these results in Theorem 4.1 below.
Theorem 4.1.
@@
(22) $\rho\left(\tilde{T}_{\star,u_{\ell}^{(1)}}\right)\leq\rho\left(\tilde{T}_{\star,u_{\ell}^{(2)}}\right),\quad\mathrm{where}\;\star\in\{\mathrm{AS},\mathrm{RAS},\mathrm{MS}\}.$
@@
(23) $\mathcal{L}u=f\quad\mathrm{in}\;\Omega\quad\mathrm{and}\quad u=g\quad\mathrm{on}\;\partial\Omega,$
@@
(24) $\mathcal{L}u:=\eta(\mathbf{x})u-\mathrm{div}\left(\alpha(\mathbf{x})\nabla u\right)+\mathbf{b}(\mathbf{x})\cdot\nabla u.$
@@
again with $\beta=100$. To discretize we use the standard 5-point stencil finite difference scheme, adapting some of the code from [26] and obtain systems of linear equations (1) with $A$ being a non-symmetric $M$-matrix. We then partition $A$ into two overlapping subdomain problems, taking the size of the overlap block to correspond to the bandwidth of $A$, i.e., we consider two overlapping subdomains $\Omega_{1},\Omega_{2}\subset\Omega$ with overlap widthAs a result, we expect the convergence factor of Schwarz method to deteriorate as $N$ increases, see [18, Section 5] and [3, Section 5]. $\mathcal{O}(h)$. We take our right-hand side vector $\mathbf{f}$ and our initial approximation vector $\mathbf{u}^{(0)}$ as random vectors with entries in $(0,1)$.
@@
As we kept the problem size relatively small so far, we next experiment also with varying $N$. However, letting $N$ grow, two numerical limitations come forward – (i) the chop toolbox becomes too slow and (ii) the verification of the condition (19) becomes untenable. Hence, for the following experiments we will use only the advanpix toolbox and only verify the condition (16) (essentially testing whether the observation in Remark 4.2 holds true also for larger $N$).
@@ 4.2 The symmetric case
First, we would like to highlight that Theorem 4.1 applies to this case “as is”. Moreover, if the rescaling and rounding is done symmetrically (or if we store and work only on, say, the upper-triangular part of the matrix), then the symmetry is preserved as well. Notably, to achieve symmetrical scaling, the proposed rescaling algorithm needs to be symmetrized, leading to an iterative procedure, see [32, Algorithms 2.5]. Moreover, our rounding routine can be further tailored to preserve other useful properties of the subdomain matrices.
@@
Next, we turn our attention to the classical convergence theory for the algebraic Schwarz methods for the symmetric, positive-definite case. As the RAS method is inherently non-symmetric, it is standard to consider the convergence theory only for the (damped) AS and MS methods and as a result we focus only on these two classes******Some theory for SPD matrices has been developed for variants of the RAS method, see [8], and recently, the convergence of RAS for SPD matrices was studied in [43] using the variational methods for a simple model problem. In general, convergence of RAS is usually addressed in combination with the particular problem, see [15], or based on other properties of the system matrix, see [18].. For these, the driving force behind Theorem 2.1 is the so-called $P$-regular Splitting Theorem, see, e.g., [38, Theorem 7.1.9]. The core assumption there becomes that the splittings in (11) are $P$-regular splitting, i.e., that
(26) $\tilde{A}_{i}^{T}+\tilde{A}_{i}-A_{i}\succeq 0.$
@@
(27) $\tilde{\mathcal{A}}_{i}\succ 0\quad\mathrm{and}\quad\lambda_{min}(\tilde{\mathcal{A}}_{i})\geq\left|\lambda_{-\infty}(\mathtt{F}_{i})\right|,$
@@
(28) $\lambda_{min}(\tilde{\mathcal{A}}_{i})\geq\lambda_{min}(\mathcal{A}_{i})+\lambda_{-\infty}(\mathtt{F}_{i}),$
@@
(29) $\lambda_{min}(\mathcal{A}_{i})\geq 2\left|\lambda_{-\infty}(\mathtt{F}_{i})\right|.$
@@
We note that analysis for SPD matrices and multiprecision additive Schwarz methods has been considered elsewhere; see [47, 2, 44, 27]. However, in all of these papers the authors consider the (non-damped) AS as a preconditioner for CG and hence the analysis and/or numerical investigation focus on the preconditioned CG method, e.g., using variational techniques that allow establishing a bound on the condition number of the preconditioned system. We also note that the assumptions look somewhat similar††††††Compare [44, equations (12) and (14)] and [2, Section 5] with (16) and (29).. Indeed, to get a coarser version of (29) we can replace $\lambda_{-\infty}(\mathtt{F}_{i})$ with $\max\limits_{\lambda\;\in\;\sigma(\mathtt{F}_{i})}|\lambda|\equiv-\rho(\mathtt{F}_{i})$ so that instead of (29) we would require
(30) $\lambda_{min}(\mathcal{A}_{i})\geq 2\rho(\mathtt{F}_{i}).$
@@
Last, we use the preconditioned GMRES with multiprecision dAS‡‡‡‡‡‡In practice, we would take advantage of the symmetry of the dAS as a preconditioner and would run a left-preconditioned CG. Here we use GMRES simply to keep the preconditioner results comparable to all of the other methods., RAS and MS as the left preconditioners and with relative residual tolerance $10^{-12}$, zero initial approximation and maximum number of iterations set to 100. We show the GMRES convergence curves and the number of iterations in Figure 10. We see that the number of iterations again stays mostly stable with respect to changing $d_{\ell}$ for a fixed $N$ and, moreover, the conditions (16) or (29) still work as a reasonably accurate indicator for the choice of the number of digits $d_{\ell}$.

### v2

@@ 1 Introduction
$A\mathbf{u}=\mathbf{f},\quad A\in\mathbb{R}^{N\times N},\mathbf{f}\in\mathbb{R}^{N}.$ (1)
@@ 2 Algebraic Schwarz methods
$\bar{R}_{i}=\begin{bmatrix}I_{\bar{N}_{i}}0\end{bmatrix}\overline{\Pi}_{i}\in\mathbb{R}^{\bar{N}_{i}\times N},$ (2)
@@
$\mathbf{u}^{(n+1)}=T_{\star}\mathbf{u}^{(n)}+\mathbf{c}_{\star},\quad\mathrm{or,\;equivalently,\;for\;the\;errors\;}\quad\mathbf{e}^{(n+1)}=T_{\star}\mathbf{e}^{(n)},$ (3)
where $\star\in\{\mathrm{AS},\mathrm{RAS},\mathrm{MS}\}$, $\mathbf{c}_{\star}\in\mathbb{R}^{N}$ are some constant vectors, $\mathbf{e}^{(n)}:=\mathbf{u}-\mathbf{u}^{(n)}\in\mathbb{R}^{N}$ is the error vector after $n$ iterations and the matrices $T_{\star}$ are the iteration matrices of the respective methods, given by [^3] An equivalent multi-splitting formulation of these can be found in [18, Section 2] and [3, Sections 2 and 3].
@@
$M_{\mathrm{MS}}^{-1}=\left(I_{N}-T_{\mathrm{MS}}\right)A^{-1},$ (5)
@@
$A^{T}=A\quad\mathrm{and}\quad\mathbf{v}^{T}A\mathbf{v}>0\;\mathrm{for\;all\;}\mathbf{v}\neq 0.$ (6)
@@
Theorem 1 ([17, Lemma 2.8][3, Theorem 3.8]).
@@
Theorem 2 ([17, Theorem 3.4][18, Theorem 4.4][3, Theorem 3.5]).
@@
Remark 1.
We note that Theorem 2 seemingly guarantees RAS convergence even for “no overlap” case where $W_{i}=\overline{W}_{i}$ for some (or all) $i$. At first this might seem contradictory to the analytic results for the standard Laplace test problem, where $W_{i}$ corresponds to discretization of the problem on $\Omega_{i}$. But there is no contradiction as even the “no overlap” case in the algebraic sense corresponds to the “$h$ overlap” in the analytic sense of the subdomains $\Omega_{i}$, thanks to the Dirichlet boundary condition enforcement.
@@ 3 Multiprecision computations
The chop toolbox is an open-source MATLAB toolbox [^4] Towards the end of preparing the manuscript, the chop toolbox has been also released for python, see [9]. developed for simulating different precisions using the native double of MATLAB, essentially by removing a portion of the mantissa of the result after each operation, corresponding to “rounding” back to the simulated precision. For computations in single precision or lower, chop faithfully simulates the computation in the precision (see [31, Section 3.1]) and can also simulate the underflow/overflow during the computation. Although this toolbox outperforms many other options (see [31, Sections 5 and 6]), it makes some computations prohibitively time-consuming, even after adapting it to sparse matrices. Although it allows for arbitrary user-defined formats (defined by the number of bits allocated to the exponent and the significand), we will restrict ourselves to the currently standard ones, summarized in Table 1 below.
@@
$\mathcal{A}_{i}\mathbf{v}_{i}=\mu\mathbf{b}_{i}$ (8)
@@
$\mathcal{A}_{i}\hat{\mathbf{v}}_{i}=\hat{\mathbf{b}}_{i},$ (9)
@@ 4 Algebraic analysis of multiprecision Schwarz methods
In this section we give analogous results to Theorems 1 and 2 when the subdomain solves $A_{i}^{-1}$ are represented using a lower-precision in some way. The purpose of the numerical experiments here is twofold – to demonstrate the theoretical results and also to build an intuition for and understanding of the multiprecision Schwarz methods. Therefore, we will use the convergence properties (such as number of iterations or the convergence factor) to compare the results with their “full precision” counterparts, as opposed to, e.g., runtimes. All of the code used to produce these is available at https://github.com/MichalOutrata/mpSchwarz but, naturally, the code assumes that both the advanpix as well as the chop toolboxes are available.
@@
$A_{i}=\tilde{A}_{i}-\left(\tilde{A}_{i}-A_{i}\right),\quad i=1,\dotsc,p\;.$ (11)
@@ 4.1 The general case
$\tilde{A}_{i}^{-1}\geq 0\quad\mathrm{and}\quad\tilde{A}_{i}^{-1}\left(\tilde{A}_{i}-A_{i}\right)=\tilde{A}_{i}^{-1}\mathtt{E}_{i}\geq 0.$ (12)
@@
$\tilde{A}_{i}=\mu^{-1}(D_{i}^{r})^{-1}\tilde{\mathcal{A}}_{i}(D_{i}^{r})^{-1}$ (13)
@@
Here we would like to recall a useful observation for diagonal re-scaling of general stationary iterative methods [^5] We came across this observation in [29, Section 17.2, below eqn. (17.3)] but this is likely not the original reference. – if the stationary iterative method is based on a splitting A=MNA=M-N such that the entries of $M$ are multiples of the corresponding entries of $A$, then the iteration matrices for $A$ and $D_{1}AD_{2}$ are similar (i.e., have the same convergence factor) for any non-singular diagonal matrices $D_{1},D_{2}$. On one hand, this shows that for $u_{\ell}=u_{w}$, the re-scaling above doesn’t affect the asymptotic convergence rate in our case. On the the other, we are clearly interested in cases where the rounding does make a difference and through this observation we see that we can expect the convergence factor to be affected by the re-scaling.
@@
$\tilde{\mathcal{A}}_{i}^{-1}\geq 0\quad\mathrm{and}\quad\tilde{\mathcal{A}}_{i}^{-1}\left(\tilde{\mathcal{A}}_{i}-\mathcal{A}_{i}\right)=\tilde{\mathcal{A}}_{i}^{-1}\mathtt{F}_{i}\geq 0.$ (14)
@@
$\tilde{\mathcal{A}}_{i}^{-1}=\mathcal{A}_{i}^{-1}\left(I+\mathtt{F}_{i}\mathcal{A}_{i}^{-1}\right)^{-1},$ (15)
@@
$\tilde{\mathcal{A}}_{i}^{-1}=\mathcal{A}_{i}^{-1}\sum\limits_{k=0}^{+\infty}\left(-\mathtt{F}_{i}\mathcal{A}_{i}^{-1}\right)^{k},$ (17)
@@
$\mathcal{A}_{i}^{-1}\geq\mathcal{A}_{i}^{-1}\mathtt{F}_{i}\mathcal{A}_{i}^{-1},$ (19)
@@
$|\mathtt{F}_{i}|\leq u_{\ell}|\mathcal{A}_{i}|,$ (20)
which is generally guaranteed (the absolute value is to be understood component-wise). Also, similarly to [2, Sections 4 and 5], both (16) and (19) invite us to choose $(u_{\ell})_{i}$ for each subdomain independently, based on the relevant quantities or their estimates [^6] In [2], the authors work with a similar idea but calculate explicitly the analogue of the inverses $\tilde{\mathcal{A}}_{i}^{-1}$ in different precisions based on their conditioning. This is somewhat complementary to our approach as our interest lies in the analysis of the resulting method rather than in the practical aspect, which has been covered in [2] and we do not comment further on how to choose $u_{\ell}$ (or $(u_{\ell})_{i}$) for the subdomain problems.. We keep (16) and (19) as assumptions, coupling the subdomain problems and the choice of the lower precision $u_{\ell}$ and move our attention to the condition $\mathtt{F}_{i}\geq 0$.
@@
To this end we assume that for any precision $u_{\ell}$ we have at our disposal the functions $\mathtt{rd}_{u_{\ell}}$ and $\mathtt{ru}_{u_{\ell}}$ that round towards zero (down) and towards plus/minus infinity (up) [^7] In the chop toolbox, these are already implemented and for the advanpix package, these are straight-forward to implement as we deal with the precision $u_{\ell}$ corresponding to $d_{\ell}$ accurate decimal digits (as opposed to dealing with bits in the case of chop).. We then introduce the rounding procedure $\mathrm{round}_{Mmtrx}()$ that for any matrix $X$ gives its low-precision approximation $\mathrm{round}_{Mmtrx}(X)$ given by
@@
$\tilde{\mathcal{A}}_{i}=\mathrm{round}_{Mmtrx}(\mathcal{A}_{i}),$ (21)
we get $\mathtt{F}_{i}\geq 0$ and obtain a convergent multiprecision Schwarz methods under the assumptions (16) and  (19); we summarize these results in Theorem 3 below.
Theorem 3.
@@
$\rho\left(\tilde{T}_{\star,u_{\ell}^{(1)}}\right)\leq\rho\left(\tilde{T}_{\star,u_{\ell}^{(2)}}\right),\quad\mathrm{where}\;\star\in\{\mathrm{AS},\mathrm{RAS},\mathrm{MS}\}.$ (22)
@@
$\mathcal{L}u=f\quad\mathrm{in}\;\Omega\quad\mathrm{and}\quad u=g\quad\mathrm{on}\;\partial\Omega,$ (23)
@@
$\mathcal{L}u:=\eta(\mathbf{x})u-\mathrm{div}\left(\alpha(\mathbf{x})\nabla u\right)+\mathbf{b}(\mathbf{x})\cdot\nabla u.$ (24)
@@
again with $\beta=100$. To discretize we use the standard 5-point stencil finite difference scheme, adapting some of the code from [26] and obtain systems of linear equations (1) with $A$ being a non-symmetric $M$-matrix. We then partition $A$ into two overlapping subdomain problems, taking the size of the overlap block to correspond to the bandwidth of $A$, i.e., we consider two overlapping subdomains $\Omega_{1},\Omega_{2}\subset\Omega$ with overlap width [^8] As a result, we expect the convergence factor of Schwarz method to deteriorate as $N$ increases, see [18, Section 5] and [3, Section 5]. $\mathcal{O}(h)$. We take our right-hand side vector $\mathbf{f}$ and our initial approximation vector $\mathbf{u}^{(0)}$ as random vectors with entries in $(0,1)$.
@@
As we kept the problem size relatively small so far, we next experiment also with varying $N$. However, letting $N$ grow, two numerical limitations come forward – (i) the chop toolbox becomes too slow and (ii) the verification of the condition (19) becomes untenable. Hence, for the following experiments we will use only the advanpix toolbox and only verify the condition (16) (essentially testing whether the observation in Remark 3 holds true also for larger $N$).
@@ 4.2 The symmetric case
First, we would like to highlight that Theorem 3 applies to this case “as is”. Moreover, if the rescaling and rounding is done symmetrically (or if we store and work only on, say, the upper-triangular part of the matrix), then the symmetry is preserved as well. Notably, to achieve symmetrical scaling, the proposed rescaling algorithm needs to be symmetrized, leading to an iterative procedure, see [32, Algorithms 2.5]. Moreover, our rounding routine can be further tailored to preserve other useful properties of the subdomain matrices.
@@
Next, we turn our attention to the classical convergence theory for the algebraic Schwarz methods for the symmetric, positive-definite case. As the RAS method is inherently non-symmetric, it is standard to consider the convergence theory only for the (damped) AS and MS methods and as a result we focus only on these two classes [^9] Some theory for SPD matrices has been developed for variants of the RAS method, see [8], and recently, the convergence of RAS for SPD matrices was studied in [43] using the variational methods for a simple model problem. In general, convergence of RAS is usually addressed in combination with the particular problem, see [15], or based on other properties of the system matrix, see [18].. For these, the driving force behind Theorem 1 is the so-called $P$-regular Splitting Theorem, see, e.g., [38, Theorem 7.1.9]. The core assumption there becomes that the splittings in (11) are $P$-regular splitting, i.e., that
$\tilde{A}_{i}^{T}+\tilde{A}_{i}-A_{i}\succeq 0.$ (26)
@@
$\tilde{\mathcal{A}}_{i}\succ 0\quad\mathrm{and}\quad\lambda_{min}(\tilde{\mathcal{A}}_{i})\geq\left|\lambda_{-\infty}(\mathtt{F}_{i})\right|,$ (27)
@@
$\lambda_{min}(\tilde{\mathcal{A}}_{i})\geq\lambda_{min}(\mathcal{A}_{i})+\lambda_{-\infty}(\mathtt{F}_{i}),$ (28)
@@
$\lambda_{min}(\mathcal{A}_{i})\geq 2\left|\lambda_{-\infty}(\mathtt{F}_{i})\right|.$ (29)
@@
We note that analysis for SPD matrices and multiprecision additive Schwarz methods has been considered elsewhere; see [47, 2, 44, 27]. However, in all of these papers the authors consider the (non-damped) AS as a preconditioner for CG and hence the analysis and/or numerical investigation focus on the preconditioned CG method, e.g., using variational techniques that allow establishing a bound on the condition number of the preconditioned system. We also note that the assumptions look somewhat similar [^10] Compare [44, equations (12) and (14)] and [2, Section 5] with (16) and (29).. Indeed, to get a coarser version of (29) we can replace $\lambda_{-\infty}(\mathtt{F}_{i})$ with $\max\limits_{\lambda\;\in\;\sigma(\mathtt{F}_{i})}|\lambda|\equiv-\rho(\mathtt{F}_{i})$ so that instead of (29) we would require
$\lambda_{min}(\mathcal{A}_{i})\geq 2\rho(\mathtt{F}_{i}).$ (30)
@@
Last, we use the preconditioned GMRES with multiprecision dAS [^11] In practice, we would take advantage of the symmetry of the dAS as a preconditioner and would run a left-preconditioned CG. Here we use GMRES simply to keep the preconditioner results comparable to all of the other methods., RAS and MS as the left preconditioners and with relative residual tolerance $10^{-12}$, zero initial approximation and maximum number of iterations set to 100. We show the GMRES convergence curves and the number of iterations in Figure 10. We see that the number of iterations again stays mostly stable with respect to changing $d_{\ell}$ for a fixed $N$ and, moreover, the conditions (16) or (29) still work as a reasonably accurate indicator for the choice of the number of digits $d_{\ell}$.